OpenStudy (minimato):

Let (-5,3) be a point on the terminal side of theta. Find the exact values of cos(theta), csc(theta), and tan (theta)

OpenStudy (evoker):

FInd r using the Pythagoras equation, then sin(theta)=y/r cos(theta)=x/r tan(theta)=y/x csc(theta)=r/y sec(theta)=r/x cot(theta)=x/y

OpenStudy (minimato):

Like a^2 + b^2 = c^2? Would r be c?

OpenStudy (minimato):

@Evoker

OpenStudy (evoker):

yeah

OpenStudy (evoker):

so for your case r is square root of 34.

OpenStudy (minimato):

Okay! And then I just substitute the square root of 34 for r into those formulas?

OpenStudy (evoker):

yep

OpenStudy (minimato):

So the answers would be cos = -5/sqrt34 csc = sqrt34/3 tan = 3/-5

OpenStudy (minimato):

Is this correct?

OpenStudy (evoker):

depends does your text want to rationalize denominators? but if not yes